# PRoLoRA: Partial Rotation Empowers More Parameter-Efficient LoRA

### v1

@@ Abstract
With the rapid scaling of large language models (LLMs), serving numerous LoRAs concurrently has become increasingly impractical, leading to unaffordable costs and necessitating more parameter-efficient finetuning methods. In this work, we introduce Partially Rotation-enhanced Low-Rank Adaptation (PRoLoRA), an intra-layer sharing mechanism comprising four essential components: broadcast reduction, rotation enhancement, partially-sharing refinement, and rectified initialization strategy. As a superset of LoRA, PRoLoRA pertains its advantages, and effectively circumvent the drawbacks of peer parameter-sharing methods with superior model capacity, practical feasibility, and broad applicability. Empirical experiments demonstrate the remarkably higher parameter efficiency of PRoLoRA in both specific parameter budget and performance target scenarios, and its scalability to larger LLMs. Notably, with one time less trainable parameters, PRoLoRA still outperforms LoRA on multiple instruction tuning datasets. Subsequently, an ablation study is conducted to validate the necessity of individual components and highlight the superiority of PRoLoRA over three potential variants. Hopefully, the conspicuously higher parameter efficiency can establish PRoLoRA as a resource-friendly alternative to LoRA.
@@
Sheng ${}^{\heartsuit}$, Boyang ${}^{\spadesuit}$, Jiacheng ${}^{\heartsuit}$, Jiyue ${}^{\spadesuit}$, Liheng ${}^{\heartsuit}$, Lingpeng ${}^{\heartsuit}$, Chuan ${}^{\heartsuit}$$^{\heartsuit}$ The University of Hong Kong, ${}^{\spadesuit}$ The Chinese University of Hong Kong u3009618@connect.hku.hk, byxue@se.cuhk.edu.hk, carsonye@connect.hku.hk, jiangjy@link.cuhk.edu.hk, clh648@connect.hku.hk, {lpk, cwu}@cs.hku.hk
@@ 1 Introduction
With the impressive capabilities, finetuning large language models (LLMs), such as LLaMA 2 (Touvron et al., 2023), GPT-3.5 Turbo (OpenAI, 2023), and Gemini (Team et al., 2023), for specific domains and functions (e.g., model alignment (Wang et al., 2023b) and instruction tuning (Zhang et al., 2023b) ), have become increasingly popular. To alleviate the high costs associated with full finetuning, parameter-efficient finetuning (PEFT), especially LoRA (Hu et al., 2021), has emerged as a lightweight solution by tuning a minority of parameters and freezing the remaining ones (Houlsby et al., 2019; Liu et al., 2022). However, with the rapid boost in model scales, the demand for further enhanced parameter efficiency becomes progressively more imperative, especially in multi-LoRA scenarios. As shown in the Table 1, the configuration in Dettmers et al. (2023) (i.e., applying LoRA with the rank of 64 to all linear layers) results in a significant number of trainable parameters. For a single LLaMA2-7B model, LoRA will have about 160 million parameters to be tuned, occupying 610MB of disk storage and GPU memory in inference. These numbers quickly escalate to about 360 million and 1.4GB for a LLaMA2-70B model. For multi-LoRA scenarios, such as personalization (Chen et al., 2023), multitasking (Huang et al., 2023) and Mixture of Experts (MoE) (Chen et al., 2024), this issue will dramatically exacerbate. Specifically, resource consumption will increase linearly with personalized customization, which will further experience a quadratic growth when coupled with multitasking. Hence, the unbearable costs in multi-LoRA scenarios do spark a demand for further improved parameter efficiency.
@@
Focusing on the above target, parameter sharing can serve as an effective approach. Generally, the rank plays a pivotal role in the performance of LoRA, and models perform better with higher ranks as listed in Table 2. Besides, given a specific trainable parameter budget, better performance means higher parameter efficiency. Hence, enhancing parameter efficiency can be transformed into appropriately increasing the rank of LoRA with the same parameter count. Although VeRA (Kopiczko et al., 2023) can be regarded as an attempt in this direction, its aggressive freezing operations result in limited model capacity and excessively high rank, leading to significant inference latency in multi-LoRA scenarios where LoRA modules are not merged into the pretrained weights. Subsequently, Tied LoRA (Renduchintala et al., 2023) alleviates these problems by allowing the inter-layer shared matrices to be trainable. However, its tying mechanism restricts its applicability to weights of different shapes, which are widely present among self-attention and MLP modules.
@@
To circumvent all the above drawbacks, we introduce a new approach called Partially Rotation-enhanced Low-Rank Adaptation (PRoLoRA). It features a parameter-sharing mechanism within the low-rank decomposition matrices, and consists of four essential components: broadcast reduction, rotation enhancement, partially-sharing refinement, and rectified initialization strategy. Specifically, we reparameterize the low-rank matrices with multiple chunks along the hidden dimension, and broadcast the first chunks to the others so that trainable parameters can be saved, or equivalently, the rank can be increased multiple times. Then a nearly cost-free rotation operation along the rank dimension is performed to differentiate the identical chunks for higher expressiveness. Besides, a minimal subset of ranks is reserved without sharing for further refined capacity. To ensure the same bounds for initialization as unshared parameters, we also rectify the vanilla Kaiming uniform distribution (He et al., 2015) for shared ones. As a superset of LoRA, PRoLoRA not only pertains the advantages of LoRA, such as lightweight task switching and optional merging to eliminate extra latency, but also brings about better capacity, practical feasibility, and broader applicability than other parameter-sharing methods. Empirical experiments on multiple instruction tuning datasets validate the higher parameter efficiency of PRoLoRA than baselines via two alternative perspectives (i.e., a specific trainable parameter budget and performance target). Impressively, with half of tunable parameters, PRoLoRA achieves 4/6 wins and better average performance over LoRA. When scaling up to LLaMA2-13B, PRoLoRA consistently outperforms LoRA with the same trainable parameter count. Additionally, comprehensive ablation studies demonstrate the necessity of each component and the superiority of PRoLoRA over three potential intra-layer sharing variants. Overall, PRoLoRA achieves significantly higher parameter efficiency, and thereby remarkably alleviates the storage and GPU memory burden in multi-LoRA scenarios, establishing PRoLoRA as a resource-friendly alternative for LoRA.
@@ 2 Related work
Parameter sharing has ever been adopted by prior studies to reduce model footprint. Universal Transformer (Dehghani et al., 2018) proposes to share all layers within a transformer model. Systematically, Takase and Kiyono (2023) refines this mechanism with three parameter-sharing strategies across transformer layers, improving the computational and parameter efficiency. Similarly, Reid et al. (2021) compares several parameter reduction methods and introduces the Subformer model with shared middle layers and embedding factorization, significantly saving the parameters without performance degradation. Subsequently, DictFormer (Lou et al., 2021) reparameterizes the original weights with a shared dictionary, unshared coefficients and indices, resulting in a more compact transformer model and faster computations. Targeting on-device deployment, EdgeFormer (Ge et al., 2022) shares the attention and FFN modules, and incorporates PEFT-based layer adaptation to minimize the number of parameters. More recently, Pires et al. (2023) removes the FFN on the decoder layers and shares a single larger FFN across the encoder, achieving substantial gains in both accuracy and latency. Differently, our study focuses on multi-LoRA scenarios and seeks to enhance the parameter efficiency of LoRA models instead of transformer models.
@@ 3.2 Advantage Analysis
As stated in Section 4.2 and 4.3, PRoLoRA achieves better performance than LoRA and others given a specific parameter budget, indicating the apparently higher parameter efficiency of PRoLoRA. In other words, for a desired performance, PRoLoRA requires less disk space during storage, fewer activation occupations during training, and lower GPU memory during inference, which significantly alleviates the burden of serving multiple models.
@@ 4.2 Main Results
Table 2: Results of LLaMA2-7B with different methods on diverse instruction following datasets. “Param.” and “Avg.” are the abbreviations of “Parameter Count” and “Average”, while the symbols ${}^{-r}$ and ${}^{-i}$ denote the ablation of the rotation enhancement and rectified initialization strategy, respectively.
@@ 4.4 Ablation Study
Figure 1: Performance of PRoLoRA with the rank of 32 with respect to unshared ranks and learning rates given a specific parameter budget on the LLaMA2-7B model and BBH benchmark.
We first study the impact of broadcast reduction and partially-sharing refinement on the performance. In Figure 1, we report the joint effects of unshared rank $u$ and learning rate. Specifically, with the unshared rank as 0, all trainable parameters are shared chunk-wisely, nullifying the partially-sharing mechanism. With the increase of the unshared rank, fewer parameters are shared, resulting in larger sharing ratios to keep the same rank (i.e., 32). Once the unshared rank reaches 8, all parameters are no longer shared, degrading PRoLoRA to vanilla LoRA. Clearly, as the sharing ratios increase with the presence of broadcast reduction, the optimal learning rate gradually increases as well. This could be attributed to the shared parameters at different locations canceling each other out during gradient updates, necessitating a larger learning rate for effective parameter optimization. This also implies that shared and unshared parameters may require different learning rates, and setting them separately could further enhance PRoLoRA’s performance, which is left for future work. However, despite the unified learning rate, PRoLoRA consistently outperforms vanilla LoRA (i.e., $u=8$) with the unshared rank spanning from 4 to 7, demonstrating the superiority of increased rank through intra-layer sharing and the parameter robustness of PRoLoRA to the unshared rank. Besides, the inferior performance without unshared parameters highlights the necessity of partial sharing refinement in PRoLoRA.
@@
Sharing / Rotaion Hidden. Rank.
@@ 6 Limitation
  2. Setting separate learning rates for shared and unshared parameters may further improve the performance of PRoLoRA. As discussed in Section 4.4, as the sharing ratios increase, larger learning rates are required to prevent mutual cancellation of updates at different positions. However, we do not apply separate learning rates in our implementation, despite the potential additional benefits for PRoLoRA. As a remedy, grouping parameters with different learning rates in the optimizer or adding independent scalars, inspired by LoRA, to the unshared parameters hold promise for extra performance enhancement of PRoLoRA and warrant investigation in future studies.
@@ 7 Ethics Statement
We strictly follow the ACL Code of Ethics during the research. To the best of our knowledge, there are no foreseeable potential risks in the methods we introduced. We report the computing infrastructure for all computational experiments presented in the paper. The transparent statistics on our results and detailed configuration of our experimental setup, including best-found hyperparameter values, are well stated. Besides, we will also release the code upon publication for publicly available reproducibility with minimal effort.

### v2

@@ Abstract
With the rapid scaling of large language models (LLMs), serving numerous low-rank adaptations (LoRAs) concurrently has become increasingly impractical, leading to unaffordable costs and necessitating more parameter-efficient finetuning methods. In this work, we introduce Partially Rotation-enhanced Low-Rank Adaptation (PRoLoRA), an intra-layer sharing mechanism comprising four essential components: broadcast reduction, rotation enhancement, partially-sharing refinement, and rectified initialization strategy. As a superset of LoRA, PRoLoRA retains its advantages, and effectively circumvent the drawbacks of peer parameter-sharing methods with superior model capacity, practical feasibility, and broad applicability. Empirical experiments demonstrate the remarkably higher parameter efficiency of PRoLoRA in both specific parameter budget and performance target scenarios, and its scalability to larger LLMs. Notably, with one time less trainable parameters, PRoLoRA still outperforms LoRA on multiple instruction tuning datasets. Subsequently, an ablation study is conducted to validate the necessity of individual components and highlight the superiority of PRoLoRA over three potential variants. Hopefully, the conspicuously higher parameter efficiency can establish PRoLoRA as a resource-friendly alternative to LoRA.
@@
Sheng Wang, Boyang Xue, Jiacheng Ye, Jiyue Jiang, Liheng Chen, Lingpeng Kong, Chuan Wu The University of Hong Kong, The Chinese University of Hong Kong u3009618@connect.hku.hk, byxue@se.cuhk.edu.hk, carsonye@connect.hku.hk, jiangjy@link.cuhk.edu.hk, clh648@connect.hku.hk, {lpk, cwu}@cs.hku.hk
@@ 1 Introduction
Finetuning large language models (LLMs), such as LLaMA 2 (Touvron et al., 2023), GPT-3.5 Turbo (OpenAI, 2023), and Gemini (Team et al., 2023), for specific domains and functions (e.g., model alignment (Wang et al., 2023b) and instruction tuning (Zhang et al., 2023b) ), have become increasingly popular. To alleviate the high costs associated with full finetuning, parameter-efficient finetuning (PEFT), especially LoRA (Hu et al., 2021), has emerged as a lightweight solution by tuning a minority of parameters and freezing the remaining ones (Houlsby et al., 2019; Liu et al., 2022). However, with the rapid boost in the number of model’s parameters, the demand for further enhanced parameter efficiency becomes progressively more imperative, especially when multiple LoRA are deployed simultaneously. As shown in the Table 1, the configuration in Dettmers et al. (2023) (i.e., applying LoRA with the rank of 64 to all linear layers) results in a significant number of trainable parameters. For a single LLaMA2-7B model, LoRA will have about 160 million parameters to be tuned, occupying 610MB of disk storage and GPU memory in inference. These numbers quickly escalate to about 360 million and 1.4GB for a LLaMA2-70B model. For multi-LoRA scenarios, such as personalization (Chen et al., 2023) and multitasking (Huang et al., 2023), this issue will dramatically exacerbate. Specifically, resource consumption will increase linearly with personalized customization, which will further experience a quadratic growth when coupled with multitasking. Hence, the high costs in multi-LoRA scenarios do spark a demand for further improved parameter efficiency.
@@
Focusing on the above target, parameter sharing can serve as an effective approach. Although small ranks could provide competitive performance in specific tasks  (Hu et al., 2021), models generally perform better with higher ranks as listed in Table 2. Besides, given a specific trainable parameter budget, better performance means higher parameter efficiency. Hence, enhancing parameter efficiency can be transformed into appropriately increasing the rank of LoRA with the same parameter count. Although VeRA (Kopiczko et al., 2023) can be regarded as an attempt in this direction, its aggressive freezing operations result in limited model capacity and excessively high rank, leading to significant inference latency in multi-LoRA scenarios where LoRA modules are not merged into the pretrained weights. Subsequently, Tied LoRA (Renduchintala et al., 2023) alleviates these problems by allowing the inter-layer shared matrices to be trainable. However, its tying mechanism restricts its applicability to weights of different shapes, which are widely present among self-attention and MLP modules.
@@
To circumvent all the above drawbacks, we introduce a new approach called Partially Rotation-enhanced Low-Rank Adaptation (PRoLoRA). It features a parameter-sharing mechanism within the low-rank decomposition matrices, and consists of four essential components: broadcast reduction, rotation enhancement, partially-sharing refinement, and rectified initialization strategy. Specifically, we reparameterize the low-rank matrices with multiple chunks along the hidden dimension, and broadcast the first chunks to the others so that trainable parameters can be saved, or equivalently, the rank can be increased multiple times. Then a nearly cost-free rotation operation along the rank dimension is performed to differentiate the identical chunks for higher expressiveness. Besides, a minimal subset of ranks is reserved without sharing for further refined capacity. To ensure the same bounds for initialization as unshared parameters, we also rectify the vanilla Kaiming uniform distribution (He et al., 2015) for shared ones. As a superset of LoRA, PRoLoRA not only pertains the advantages of LoRA, such as lightweight task switching and optional merging to eliminate extra latency, but also brings about better capacity, practical feasibility, and broader applicability than other parameter-sharing methods. Empirical experiments on multiple instruction tuning datasets validate the higher parameter efficiency of PRoLoRA than baselines via two alternative perspectives (i.e., a specific trainable parameter budget and performance target). With half of tunable parameters, PRoLoRA achieves 4/6 wins and better average performance over LoRA. When scaling up to LLaMA2-13B, PRoLoRA consistently outperforms LoRA with the same trainable parameter count. Additionally, comprehensive ablation studies demonstrate the necessity of each component and the superiority of PRoLoRA over three potential intra-layer sharing variants. Overall, PRoLoRA achieves significantly higher parameter efficiency, and thereby remarkably alleviates the storage and GPU memory burden in multi-LoRA scenarios, establishing PRoLoRA as a resource-friendly alternative for LoRA.
@@ 2 Related work
Parameter sharing has been adopted by prior studies to reduce model footprint. Universal Transformer (Dehghani et al., 2018) proposes to share all layers within a transformer model. Systematically, Takase and Kiyono (2023) refines this mechanism with three parameter-sharing strategies across transformer layers, improving the computational and parameter efficiency. Similarly, Reid et al. (2021) compares several parameter reduction methods and introduces the Subformer model with shared middle layers and embedding factorization, significantly saving the parameters without performance degradation. Subsequently, DictFormer (Lou et al., 2021) reparameterizes the original weights with a shared dictionary, unshared coefficients and indices, resulting in a more compact transformer model and faster computations. Targeting on-device deployment, EdgeFormer (Ge et al., 2022) shares the attention and FFN modules, and incorporates PEFT-based layer adaptation to minimize the number of parameters. More recently, Pires et al. (2023) removes the FFN on the decoder layers and shares a single larger FFN across the encoder, achieving substantial gains in both accuracy and latency. Differently, our study focuses on multi-LoRA scenarios and seeks to enhance the parameter efficiency of LoRA models instead of transformer models.
@@ 3.2 Advantage Analysis
As stated in Section 4.2 and 4.3, PRoLoRA achieves better performance than LoRA and others given a specific parameter budget, indicating the apparently higher parameter efficiency of PRoLoRA. In other words, for a desired performance, PRoLoRA requires less disk space during storage, and lower GPU memory during inference, which significantly alleviates the burden of serving multiple models.
@@ 4.2 Main Results
Table 2: Results of LLaMA2-7B with different methods on diverse instruction following datasets. “Param.” and “Avg.” are the abbreviations of “Parameter Count” and “Average”, while the symbols -r and -i denote the ablation of the rotation enhancement and rectified initialization strategy, respectively. “4/8” and “16/32” means raising the rank to either 4 or 8, and 16 or 32, respectively. Underlined represents the optional higher ranks, while bold indicates the best result for each benchmark.
@@ 4.4 Ablation Study
Figure 4: Performance of PRoLoRA with the rank of 32 with respect to unshared ranks and learning rates given a specific parameter budget on the LLaMA2-7B model and BBH benchmark. Specially, when the unshared rank is 8, all the ranks are unshared (i.e., vanilla LoRA).
We first study the impact of broadcast reduction and partially-sharing refinement on the performance. In Figure 4, we report the joint effects of unshared rank $u$ and learning rate. Specifically, with the unshared rank as 0, all trainable parameters are shared chunk-wisely, nullifying the partially-sharing mechanism. With the increase of the unshared rank, fewer parameters are shared, resulting in larger sharing ratios to keep the same rank (i.e., 32). Once the unshared rank reaches 8, all parameters are no longer shared, degrading PRoLoRA to vanilla LoRA. Clearly, as the sharing ratios increase with the presence of broadcast reduction, the optimal learning rate gradually increases as well. This implies that shared and unshared parameters may require different learning rates, and setting them separately could further enhance PRoLoRA’s performance, which is left for future work. However, despite the unified learning rate, PRoLoRA consistently outperforms vanilla LoRA (i.e., $u=8$) with the unshared rank spanning from 4 to 7, demonstrating the superiority of increased rank through intra-layer sharing and the parameter robustness of PRoLoRA to the unshared rank. Besides, the inferior performance without unshared parameters highlights the necessity of partial sharing refinement in PRoLoRA.
@@
Sharing / Rotation Hidden. Rank.
@@ 6 Limitation
  2. As discussed in Section 4.4, setting separate learning rates for shared and unshared parameters may further improve the performance of PRoLoRA. However, we do not apply separate learning rates in our implementation, despite the potential additional benefits for PRoLoRA. As a remedy, grouping parameters with different learning rates in the optimizer or adding independent scalars, inspired by LoRA, to the unshared parameters hold promise for extra performance enhancement of PRoLoRA and warrant investigation in future studies.
  3. We claim that PRoLoRA requires lower GPU memory during inference in Section 3.2, even if our implementation of broadcasting allocates new memory to store the copied chunks, thereby negating any memory reduction for one module. However, we argue that it does not imply that overall memory usage cannot be reduced. Specifically, there are lots of PRoLoRA matrices in a model, and they are not computed simultaneously. When computing the preceding modules, the subsequent PRoLoRA matrices do not need to be copied. Once the computation of the preceding modules is completed, their occupied memory can be released promptly. Compared to the doubled parameter efficiency, we believe that the memory occupied by temporary copies can be negligible. In multi-user scenarios, only the active PRoLoRA matrices for active users need to be expanded, while other PRoLoRA matrices can still save lots of memory. Besides, we also call for the efficient implementation of PRoLoRA with CUTLASS  [^3] https://github.com/NVIDIA/cutlass and other relevant techniques.
@@ 7 Ethics Statement
We strictly follow the ACL Code of Ethics during the research. To the best of our knowledge, there are no foreseeable potential risks in the methods we introduced. We report the computing infrastructure for all computational experiments presented in the paper. The transparent statistics on our results and detailed configuration of our experimental setup, including best-found hyperparameter values, are well stated. Besides, we will also release the code upon publication for publicly available reproducibility with minimal effort. test (Brown et al., 2020)
8 Acknowledgement
This work was supported in part by Hong Kong Innovation and Technology Support Programme Platform Research Project fund (ITS/269/22FP), the joint research scheme of the National Natural Science Foundation of China (NSFC) and Hong Kong Research Grants Council (RGC) (under grant N_HKU714/21), and RGC grants 17204423 and C7004-22G (CRF).